# SplatSim: Zero-Shot Sim2Real Transfer of RGB Manipulation Policies Using Gaussian Splatting

### v2

@@ II-A Sim2real
Robotics simulation tools like [10, 11, 12, 13, 14, 15] have become invaluable in scaling up robot learning due to several advantages including parallelization, cost and time efficiency, and safety. Recent advancements in transferring learned policies from simulation to the real world have demonstrated impressive results, particularly in domains that leverage modalities with a low Sim2Real gap, such as depth, point cloud, proprioception, or tactile feedback. These modalities have enabled robots to perform contact-rich tasks like quadruped locomotion [1, 16, 2, 17, 18] and bipedal locomotion [19, 20], dexterous manipulation [3, 4, 5, 6, 7], manipulation of articulated objects [21, 22], among others [23, 24, 25, 26].
However, Sim2Real transfer for RGB-based manipulation policies remains challenging. Previous attempts in this domain have primarily focused on tasks like navigation [27], where high-fidelity collision meshes from scans [28, 29] can enhance visual realism but fail to capture dynamic interactions with the environment. Furthermore, existing approaches such as domain adaptation [30], often rely on extensive offline data collection of real-world object interactions.
In this work, we address the challenge of transferring RGB-based policies for manipulation tasks, which require rendering complex interactions between objects and the robot. Our method requires only an initial video of the static scene without the need for additional real-world data collection. A work notably related to ours is RialTo [31] which uses a Real2Sim2Real approach similar to ours. However, their policy is still trained on point clouds, which requires depth during execution time. In contrast, SplatSim only uses RGB images for learning and policy deployment. Another recent work Maniwhere [32] does large-scale reinforcement learning in simulation and shows generalization to the real world, however, their method still requires depth at test time and cannot work with just RGB images in the real world.

### v3

@@ II-A Sim2real
Robotics simulation tools like [10, 11, 12, 13, 14, 15, 16, 11] have become invaluable in scaling up robot learning due to several advantages including parallelization, cost and time efficiency, and safety. Recent advancements in transferring learned policies from simulation to the real world have demonstrated impressive results, particularly in domains that leverage modalities with a low Sim2Real gap, such as depth, point cloud, proprioception, or tactile feedback. These modalities have enabled robots to perform contact-rich tasks like quadruped locomotion [1, 17, 2, 18, 19] and bipedal locomotion [20, 21], dexterous manipulation [3, 4, 5, 6, 7], manipulation of articulated objects [22, 23], among others [24, 25, 26, 27].
However, Sim2Real transfer for RGB-based manipulation policies remains challenging. Several prior works [28, 29, 30, 31] have explored Sim2Real transfer for RGB-based manipulation policies using domain randomization techniques. These approaches, however, often require task-specific tuning and environment engineering, which can be both labor-intensive and difficult to achieve accurately in traditional physics simulators. In contrast, our method eliminates the need for task-specific adjustments and leverages an initial scan of the target deployment environment, significantly simplifying the process. Other existing approaches such as domain adaptation [32], often rely on extensive offline data collection of real-world object interactions. Our method requires only an initial video of the static scene without the need for additional real-world data collection.
In this work, we address the challenge of transferring RGB-based policies for manipulation tasks using Gaussian Splatting, which requires rendering complex interactions between objects and the robot. A work notably related to ours is RialTo [33] which uses a Real2Sim2Real approach similar to ours. However, their policy is still trained on point clouds, which requires depth during execution time. In contrast, SplatSim only uses RGB images for learning and policy deployment. Another recent work Maniwhere [34] does large-scale reinforcement learning in simulation and shows generalization to the real world, however, their method still requires depth at test time and cannot work with just RGB images in the real world.